OpenStudy (anonymous):

Write a rule for the function represented by the table. Input (x) 0 1 2 3 4 Output (y) 3 6 9 12 15

OpenStudy (anonymous):

y(0) = 3 y(1) = 6 = 3 + 3*1 y(2) = 9 = 3 + 3*2 y(3) = 12 = 3 + 3*3 So the function must be y = 3x + 3

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

can you help me with 1 more

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Identify the function rule shown in the table. n 3 4 5 6 y 2 1 0 –1

OpenStudy (anonymous):

a.y = 2 + n b.y = 5n c.y = 5 – n d.not enough information

OpenStudy (anonymous):

The answer is c.

OpenStudy (anonymous):

thx